Solve the following bvp by eigenfunction expansion method.

y''+2y=2-4x, y(0)=0, y(1)+y'(1)=0


1
Expert's answer
2021-09-13T16:47:02-0400

Characteristics equation:

"k^2+2=0\\roots\space are\\
k_{1,2}=\pm\sqrt{2}\\"

Basic solutions:

"y_2(x)=cos(\sqrt{2}\cdot x)\\
y_1(x)=sin(\sqrt{2}\cdot x)\\"

General solution of homogeneous part:

"y_h(x,C_1,C_2)=C_1\cdot cos(\sqrt{2}\cdot x)+
C_2\cdot sin(\sqrt{2}\cdot x);"

Let find partial solution in form:

"y_1=a+b\cdot x"

For finding a,b substitute it in diff equation:

"y_1''+2\cdot y_1=2a+2bx=2-4x"

We take a,b values:

a=1, b=-2;

Partial solution has the form:

"y_1(x)=1-2\cdot x"

The overall solution as a whole

"y_{nh}(x,C_1,C_2)=C_1\cdot cos(\sqrt{2}\cdot x)+

C_2\cdot sin(\sqrt{2}\cdot x)+1-2\cdot x"

Now we apply boundary condition y(0)=0 and have:

0="C_1+1"

From it we have "C_1=-1"

And therefore

"y(x,C_2)= -cos(\sqrt{2}\cdot x)+

C_2\cdot sin(\sqrt{2}\cdot x)+1-2\cdot x"

Further we use second boundary condition:

"y(1)+y'(1)=-cos(\sqrt 2)+C_2\cdot sin(\sqrt 2)-1+\sqrt 2\cdot \\

sin(\sqrt 2)+\sqrt 2\cdot C_2\cdot cos(\sqrt 2)-2=0"

This is an equation with respect to C2

Solving it we have:

"C_2=\frac{3+cos(\sqrt 2)-\sqrt 2\cdot sin(\sqrt 2)}{sin(\sqrt 2)+\sqrt 2\cdot cos(\sqrt 2)}"

And finally

"y(x,)= -cos(\sqrt{2}\cdot x)+\\
\left( \frac{3+cos(\sqrt 2)+\sqrt 2\cdot sin(\sqrt 2)}{sin(\sqrt 2)+\sqrt 2\cdot cos(\sqrt 2)} \right )\cdot sin(\sqrt{2}\cdot x)+1-2\cdot x"

Rhus we have sold bde problem.